OpenStudy (anonymous):

How do I solve this?? Choose the correct function value for: cot 81.1°

OpenStudy (jdoe0001):

and the choices to "choose" from are?

OpenStudy (anonymous):

a) 0.1566 b) 0.8434 c) 0.1723 d) 0.6826

OpenStudy (anonymous):

\[\cot \left( 81.9 \right)=\cot \left( 90-8.9 \right)=\tan 8.9\] i think now you can solve it.

OpenStudy (anonymous):

so.. plug tan8.9 into my calculator?

OpenStudy (anonymous):

okay, i figured it out